OpenStudy (anonymous):

find the 4 term of the sequence (recursive) a1= 12 an=1/2 an-1

zepdrix (zepdrix):

\[\Large\rm a_1=12\]\[\Large\rm a_n=\frac{1}{2}a_{n-1}\] So it looks like each term will be half as large as the previous term.\[\Large\rm a_2=\frac{1}{2}a_{2-1}\]\[\Large\rm a_2=\frac{1}{2}a_{1}\]\[\Large\rm a_2=\frac{1}{2}\cdot 12\] \[\Large\rm a_2=6\]

zepdrix (zepdrix):

Understand how that process works? :o Just cut that number in half a couple more times in order to find the `fourth` term.